An investment of $4500 is made at an annual simple interest rate of 9.25%. How much additional money must be invested at an annual simple interest rate of 13.5% so that the total interest earned is 12% of the total investment?
Answer by mananth(16946)

$4500--------------9.25%
x............13.5%
Total yield = 12%
4500*9.5%+13.5%(x)=12%*(4500+x)
multiply by 100
4500*9.5+13.5(x)=12(4500+x)
42750+13.5x=154000+12x
13.5x-12x=54000-42750
1.5x=11250
x= 7500

$ 7500 has to be invested to get a total return of 12%